OpenStudy (anonymous):

i dont know how to solve this=p 81x^2-49=0 the next step is 81x2-49=0 right? what do i do next?

OpenStudy (anonymous):

add \(49\) to both sides, you will get \[81x^{2}=49\] then take the square root of both sides, \[\pm9x=\pm7\] then divide by \(\pm9\) on both sides, \[x=\pm\frac{7}{9}\]

OpenStudy (anonymous):

i havent learned that yet..

OpenStudy (anonymous):

which part? I'll explain to you slower.

OpenStudy (anonymous):

anyone else?

OpenStudy (anonymous):

\[81x^{2}-49+49=0+49\] do you get this part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{81x^2}=\sqrt{49}\] do you get this part?

OpenStudy (anonymous):

not really whya re you doing that?

OpenStudy (anonymous):

because the square root of the left hand side will always equal to the square root of the right hand side. for example. \[\sqrt{9}=\sqrt{3*3}\]or\[\sqrt{4}=\sqrt{\frac{8}{2}}\]

OpenStudy (anonymous):

do you understand this part? i will explain in detail if you dont :)

OpenStudy (raden_zaikaria):